Project

General

Profile

Actions

2015-06-05 Physics coordinator report

Step IV

  • Physics shifts
    • Have filled 2015/01a
    • 2015/01b is empty - it should be fun though
  • Beam optics
    • Jaroslaw has provided optics settings for 140, 200, 240 settings;
  • Magnet misalignment
    • factored in John Cobb estimated alignments of "2015-05-11", tracking still looks okay
    • seek to look at the actual installed alignments (waiting for data to come through the system)

Step I

  • EMR paper
    • Version two of the MICE note has gone to referees
  • Pion contamination paper
    • Revised analysis is complete, MICE note in preparation

MICE Demonstration of Ionisation Cooling

  • Paper is in circulation among authors

Other

  • CM42 proto-agenda distributed to mice-analysis list - will fill indico early next week

Updated by Rogers, Chris about 8 years ago ยท 4 revisions